OpenStudy (anonymous):

Find a quadratic equation with roots –1 + 4i and –1 – 4i.

OpenStudy (anonymous):

equation in factored form will give you the roots: \[x_1 = -1+4i\] \[x_2 = -1-4i\] \[(x-x_1)(x-x_2) = 0\] substitute the values of x into the equation and expand